# Quantum amplitude estimation from classical signal processing

### v1

@@ IV.2 Parallelizability
Our algorithm is fully parallelizable, because the quantum circuits at which we take samples from is known beforehand. Here we look at the parallel query complexity for a given number of quantum computers. The idea is to divide the total number of queries as evenly as possible among the QPUs. The parallel query complexity is then simply the largest query complexity among the QPUs. We can find this even division using a greedy approach: sort the circuits from which we take samples in decreasing Grover power and distribute the circuits evenly among the QPUs starting with the highest Grover power. In Fig. 5 we plot the constant factor $C=N\cdot\varepsilon$ for the parallel query complexity given a number of QPUs. One can see that the constant factor saturates at some point (12 QPUs in this case), because the largest Grover power has to be run on some QPU, hence providing a lower bound on the parallel query complexity.
@@
In quantum computer architectures where different Grover powers can be executed on the same QPU, the parallelizability of AE also allows for the application of the technique known as Hamming weight phasing (HWP) which efficiently applies repeated arbitrary single-qubit rotations by the same angle [34, 35]. When $m$ repeated rotations by the same angle $R_{z}(\theta)$ are parallelizable, this technique reduces the number of rotations which need to be synthesized to $\lfloor\log_{2}m+1\rfloor$, at the cost of $m-1$ Toffoli gates and $m-1$ ancilla qubits. Because the Grover operator consists of repeated applications of a unitary $U$ (and $U^{\dagger}$), any single-qubit rotations in $U$ can therefore be applied using HWP across different powers of the Grover operator which have been parallelized on the QPU. The practical benefit from this approach depends on the specific form of $U$ and the total number of Grover powers that must be applied, but we highlight this ability to reduce the total number of single-qubit rotations across different Grover powers as this is not possible in iterative variants of AE.
IV.3 Noise Resilience
Here we study the ability of our approach to handle noise in the quantum circuit. This could arise from current quantum devices that do not have error correction capabilities or future fault-tolerant devices where one might be interested in understanding what size quantum error correcting code is required for a given application.
We study this by implementing a simple noise model. When estimating the probability amplitudes at a given query depth $n$ as defined in Eqs. (3) and (4), we shift the probability by a per-oracle noise parameter $\eta$. That is, for the $Z$ basis measurements, we shift the probability to
$p_{0}(n)$ $\to\eta_{n}p_{0}(n)+(1-\eta_{n})/2$ (18a)
$p_{1}(n)$ $\to\eta_{n}p_{1}(n)+(1-\eta_{n})/2,$ (18b)
where $\eta_{n}=(1-\eta)^{n}$ is the noise parameter at a given depth $n$. As the circuit depth gets deeper, the noise parameter $\eta_{n}$ gets smaller resulting in the probabilities tending towards the completely mixed state. We implement a similar model for the $X$ basis measurements.
Figure 6: Estimation error as a function of total query complexity for amplitude $a=0.5$ and confidence $\delta=0.95$ for given values of the per-oracle noise parameter $\eta$. As $\eta$ increases, the ability of the algorithm to estimate the amplitude saturates as one would expect.
Table 2: Total number of queries, maximum coherent query depth, and array parameters required to achieve a target estimation error of $\varepsilon_{\delta}=1\times 10^{-3}$ for a given noise parameter $\eta$ and confidence level $\delta$. A dashed line indicates that the noise was too high and the estimation error could not be brought below the target threshold. The results suggest that for optimal performance at this estimation error, one should target a per-oracle error rate of $\eta\leq 10^{-5}$, where one essentially achieves noise-free performance.
$\eta=10^{-3}$- -- $89,453$ $4,374$ $6,807$ $512$
$\eta=10^{-4}$ $18,262$ $1,152$ $8,399$ $648$ $2,311$ $192$
$\eta\leq 10^{-5}$ $10,214$ $512$ $6,004$ $360$ $2,311$ $192$
We plot the estimation error at $\delta=0.95$ confidence level for $a=0.5$ in Fig. 6 for a range of $\eta$ values. One can see that as the noise level increases, the ability of the algorithm to estimate the amplitude decreases exactly as one would expect. To quantify the performance of the algorithm, we report the number of queries required to achieve a given confidence level with estimation error $\varepsilon_{\delta}=1\times 10^{-3}$ which is generally consistent with many algorithmic performance requirements. These values are reported in Tab. 2.
In addition to the query counts, we also found that varying the array yielded better results. The reason, can be seen by examining Fig. 6. The target estimation error of $\varepsilon_{\delta}=1\times 10^{-3}$ lies between two arrays when using the schedule that increases the array size by powers of 2. However, as shown in Theorem II.2 the $N_{k}$ values do not need to be equal to the value 2, which we chose when reporting the scaling parameter. By choosing other integers, one can obtain an array that lies in the midpoint near where the estimation error crosses the target amount.
We tuned the array parameter values and the number of shots at each depth until we found an array that achieved the target estimation error and confidence level with the fewest total queries. This tuning was manual, so it is likely one could find better parameter values with more work. We report the total and parallel query complexities Tab. 2. We report a detailed summary of the array parameters, query depths, and shot scheduled used for those arrays in App. A.
@@ V Discussion and Future Work
We provide numerical results demonstrating one particular DOA algorithm ESPRIT that is computationally efficient, and achieves what is to our knowledge the best known query complexity by a small factor and the best known parallel query complexity by over an order of magnitude over previous approaches.
@@
We can quantify the practical impact of the AE algorithm introduced in this manuscript by considering the problem of financial derivative pricing [3, 4, 5, 6]. In Ref. [36] the authors calculate that quantum advantage in pricing an autocallable derivative contract to accuracy $2\times 10^{-3}$ with confidence $68\%$ requires a T-depth of $4.5\times 10^{7}$, T-count of $2.4\times 10^{9}$ and in order for the calculation to meet the classical target of $1$ second, a $45$ MHz logical clock rate. This calculation assumes the IQAE variant [14] is employed at $\varepsilon=10^{-3}$ and confidence $\delta=68\%$ such that the total number of oracle calls is $C/\varepsilon=5735$ [5]. From the third column of Table 1, we observe that the total number of oracle calls to achieve the same accuracy and confidence using csAE is $2000$, a 2.9$\sim$ 2.9$\times$ 2.9 × reduction in both T-depth and T-count. Moreover, because csAE allows for parallelization, assuming it can be fully harnessed across multiple QPUs, the deepest circuit requires $162$ oracle calls (sixth column of Table 1), such that the logical clock rate required for quantum advantage is in theory reduced from $45$ MHz to $1.3$ MHz, a factor of $35$x reduction. However, while this figure might be of consequence in a discussion of whether a QPU with the calculated specifications for advantage can be constructed, it assumes that we parallelize the quantum algorithm, but not the classical counterpart used for comparison, Monte Carlo. Because Monte Carlo can be parallelized in a straightforward manner, a fair comparison would allow for both quantum and classical methods to execute in a parallel fashion. From Fig. 5, we observe that the csAE runtime scales as $\sim 1/N$ for $N$ processors, precisely the same way a classical Monte Carlo simulation scales with the number of processors. This means that no further logical clock rate benefit arises from the parallelization of csAE if we allow the same parallelization to Monte Carlo. Nevertheless, csAE does provide the possibility of parallelization, enabling trade-offs between circuit width and depth when desirable, unlike other high-performing AE variants like IQAE and chebAE, overcoming an important limitation of AE methods when compared to classical Monte Carlo.
@@
There are a number of avenues to further improve or generalize our results depending on specific use cases. DOA estimation algorithms typically assume that the noise contained in the estimated signal is Gaussian and uncorrelated. However, the estimator we use to convert the quantum measurements to the form needed for DOA estimated, given in Eq. (5), has noise arising from the ratio of two binomially-distributed noise terms, which is manifestly non-Gaussian. In addition the estimator is biased. Despite this, our numerical results using an “off the shelf” implementation of ESPRIT already beat the previous best published results for worst-case complexity. We leave for future work further optimizations that may lead to even better constant factors by taking into account the actual noise distribution. Results in the DOA literature suggest that this can lead to substantial improvements (see e.g. Ref. [37]). We further remark that the noise resilience of this approach may make this method well-suited to NISQ-based algorithms that require amplitude-estimation as well [38, 39].
@@
We chose a particular sampling strategy that we refer to as the $2q$-array. This choice, however, is not limiting and in fact there is a great degree of flexibility. We demonstrated the one degree of freedom that $2q$ arrays have in that the schedule can be fine-tuned to meet a required error tolerance, as shown in Sec. IV.3. One can also choose different array geometries such as co-prime arrays [24], ruler arrays [40], fractal arrays [41], along with many other approaches [42, 43]. Different array choices can lead to different asymptotic complexity $\mathcal{O}(1/\varepsilon^{\alpha})$ for various values of $\alpha$ allowing one to achieve lower circuit depth at the cost of needing more samples. This can be useful for noisy processors where circuit depth is limited [38, 39, 16].
@@ Acknowledgements.
We thank Peter Johnson for constructive feedback on this manuscript and Patrick Rall, Bryce Fuller and Stefan Woerner for discussions on amplitude estimation.
@@ Appendix A Array Parameters
Simulation parameters $\eta=10^{-3}$; $K=1.5$
Array Parameters --
Query Depths --
Shots per Depth --
Total Queries --
Simulation parameters $\eta=10^{-4}$; $K=1.8$
Array Parameters $[3,3,2,2,2,2,2,2,2,2]$
Query Depths $[0,1,2,3,6,9,18,36,72,144,288,576,1152]$
Shots per Depth $[24,22,20,18,17,15,13,11,9,8,6,4,2]$
Total queries $18,262$
Simulation parameters $\eta=10^{-5}$; $K=2.1$
Array Parameters $[2,2,2,2,2,2,2,2,2,2]$
@@
Total Queries $10,214$
Table 4: Parameter values for $\varepsilon_{0.95}\leq 10^{-3}$.
Simulation parameters $\eta=10^{-3}$; $K=1.1$
@@
Query Depths $[0,1,2,3,6,9,18,27,54,81,162,243,486,729,1458,2187,4374]$
Shots per Depth $[31,29,27,26,24,22,20,18,17,15,13,11,9,8,6,4,2]$
Total Queries $89,453$
Simulation parameters $\eta=10^{-4}$; $K=1.1$
Array Parameters $[3,3,3,3,2,2,2,2]$
Query Depths $[0,1,2,3,6,9,18,27,54,81,162,324,648]$
Shots per Depth $[15,14,13,11,10,9,8,7,6,5,4,3,2]$
Total Queries $8,399$
Simulation parameters $\eta=10^{-5}$; $K=1.3$
Array Parameters $[6,5,3,2,2,2]$
Query Depths $[0,1,2,3,4,5,6,12,18,24,30,60,90,180,360]$
Shots per Depth $[20,19,17,16,15,13,12,11,10,8,7,6,4,3,2]$
Total Queries $6,004$
Table 5: Parameter values for $\varepsilon_{0.68}\leq 10^{-3}$.
Simulation parameters $\eta=10^{-3}$; $K=1.5$
Array Parameters $[2,2,2,2,2,2,2,2,2,2]$
Query Depths $[0,1,2,4,8,16,32,64,128,256,512]$
Shots per Depth $[17,15,14,12,11,9,8,6,5,3,2]$
Total Queries $6,807$
Simulation parameters $\eta=10^{-4}$; $K=1.1$
Array Parameters $[3,2,2,2,2,2,2,2]$
Query Depths $[0,1,2,3,6,12,24,48,96,192]$
Shots per Depth $[11,10,9,8,7,6,5,4,3,2]$
Total Queries $2,311$
Simulation parameters $\eta=10^{-5}$; $K=1.1$
Array Parameters $[3,2,2,2,2,2,2,2]$
Query Depths $[0,1,2,3,6,12,24,48,96,192]$
Shots per Depth $[11,10,9,8,7,6,5,4,3,2]$
Total Queries $2,311$

### v2

@@ IV.3 Parallelizability
The csAE algorithm is highly parallelizable, because the quantum circuits at which we take samples from is known beforehand. The results shown before provide the minimum depth circuit one can run assuming full parallelizability. Here we look to answer the question of how to parallelize the algorithm over a fixed number of QPUs. The best approach is to divide the total number of queries as evenly as possible among the available QPUs. The parallel query complexity is then simply the largest query complexity among the QPUs. We can find this even division using a greedy approach: sort the circuits from which we take samples in decreasing Grover power and distribute the circuits evenly among the QPUs starting with the highest Grover power.
In Fig. 4 we plot the maximum number of sequential applications of the circuit $U$ (which is twice the number of applications of the Grover oracles) given a number of QPUs assuming classical sampling, chebAE, and csAE. One can see that the parallelizability saturates for csAE, because the largest Grover power has to be run on some QPU, hence providing a lower bound on the parallel query complexity. Our results are presented assuming a target $\epsilon_{0.95}=10^{-3}$ consistent with row two of Tab. 2.
@@
In quantum computer architectures, where different Grover powers can be executed on the same QPU, the parallelizability of csAE also allows for the application of the technique known as Hamming weight phasing (HWP) which efficiently applies repeated arbitrary single-qubit rotations by the same angle [34, 35]. When $m$ repeated rotations by the same angle $R_{z}(\theta)$ are parallelizable, this technique reduces the number of rotations which need to be synthesized to $\lfloor\log_{2}m+1\rfloor$, at the cost of $m-1$ Toffoli gates and $m-1$ ancilla qubits. Because the Grover operator consists of repeated applications of a unitary $U$ (and $U^{\dagger}$), any single-qubit rotations in $U$ can therefore be applied using HWP across different powers of the Grover operator which have been parallelized on the QPU. The practical benefit from this approach depends on the specific form of $U$ and the total number of Grover powers that must be applied, but we highlight this ability to reduce the total number of single-qubit rotations across different Grover powers as this is not possible in iterative variants of AE.
@@ V Discussion and Future Work
We provide numerical results demonstrating one particular DOA algorithm ESPRIT that is computationally efficient which is, to our knowledge, on par with the best known query complexity and the best known parallel query complexity by over an order of magnitude over previous approaches.
@@
We can quantify the practical impact of the AE algorithm introduced in this manuscript by considering the problem of financial derivative pricing [3, 4, 5, 6]. In Ref. [36] the authors calculate that quantum advantage in pricing an autocallable derivative contract to accuracy $2\times 10^{-3}$ with confidence $68\%$ requires a T-depth of $4.5\times 10^{7}$, T-count of $2.4\times 10^{9}$ and in order for the calculation to meet the classical target of $1$ second, a $45$ MHz logical clock rate. This calculation assumes the IQAE variant [14] is employed at $\varepsilon=10^{-3}$ and confidence $\delta=68\%$ such that the total number of oracle calls is $C/\varepsilon=5735$ [5]. From the fifth column of Table 2, we observe that the total number of oracle calls to achieve the same accuracy and confidence using csAE is $1670$, a 3.43$\sim$ 3.43$\times$ 3.43 × reduction in both T-depth and T-count. Moreover, because csAE allows for parallelization, assuming it can be fully harnessed across multiple QPUs, the deepest circuit requires $102$ oracle calls (sixth column of Table 2), such that the logical clock rate required for quantum advantage is in theory reduced from $45$ MHz to $0.8$ MHz, a factor of $56$x reduction. However, while this figure might be of consequence in a discussion of whether a QPU with the calculated specifications for advantage can be constructed, it assumes that we parallelize the quantum algorithm, but not the classical counterpart used for comparison, Monte Carlo. Because Monte Carlo can be parallelized in a straightforward manner, a fair comparison would allow for both quantum and classical methods to execute in a parallel fashion. From Fig. 4, we observe that the csAE runtime scales as $\sim 1/N$ for $N$ processors, precisely the same way a classical Monte Carlo simulation scales with the number of processors. This means that no further logical clock rate benefit arises from the parallelization of csAE if we allow the same parallelization to Monte Carlo. Nevertheless, csAE does provide the possibility of parallelization, enabling trade-offs between circuit width and depth when desirable, unlike other high-performing AE variants like IQAE and chebAE, overcoming an important limitation of AE methods when compared to classical Monte Carlo.
@@
There are a number of avenues to further improve or generalize our results depending on specific use cases. One obvious example is to use more sophisticated DOA algorithms on the signal. Despite this, our numerical results using an “off the shelf” implementation of ESPRIT already is on par with the previous best published results for worst-case complexity. We leave for future work further optimizations that may lead to even better constant factors by taking into account the actual noise distribution. Results in the DOA literature suggest that this can lead to substantial improvements (see e.g. Ref. [37]). We further remark that the noise resilience of this approach may make this method well-suited to NISQ-based algorithms that require amplitude-estimation as well [38, 39].
@@
We chose a particular sampling strategy that we refer to as the $2q$-array. This choice, however, is not limiting and in fact there is a great degree of flexibility. We demonstrated the use of one particular $2q$-array, but note that there is a freedom here that one can exploit to further optimize the algorithm. One can also choose different array geometries such as co-prime arrays [24], ruler arrays [40], fractal arrays [41], along with many other approaches [42, 43]. Different array choices can lead to different asymptotic complexity $\mathcal{O}(1/\varepsilon^{\alpha})$ for various values of $\alpha$ allowing one to achieve lower circuit depth at the cost of needing more samples. This can be useful for noisy processors where circuit depth is limited [38, 39, 16].
@@ Acknowledgements.
We thank Peter Johnson for constructive feedback on this manuscript and Patrick Rall, Bryce Fuller and Stefan Woerner for discussions on amplitude estimation. We also thank Ethan Davies and Pei Zeng for finding a (serious) mistake in the previous version of the paper.
@@ Appendix A Array Parameters
Simulation parameters $K=8.1$
Array Parameters $[2,2,2,2,2,2,2,2,2]$
Query Depths $[0,1,2,4,8,16,32,64,128,256]$
Shots per Depth $[162,73,65,57,49,41,33,25,17,9]$
Total Queries $8,777$
Max Depth $256$
Table 4: Parameter values for $\varepsilon_{0.95}\leq 10^{-3}$.
Simulation parameters $K=4$
Array Parameters $[2,2,4,2,2,2,2,2]$
@@
Total Queries $4,488$
Max Depth $256$
Table 5: Parameter values for $\varepsilon_{0.68}\leq 10^{-3}$.
Simulation parameters $K=3$
@@
Query Depths $[0,1,2,4,8,16,32,64,128]$
Shots per Depth $[54,24,21,18,15,12,9,6,3]$
Total Queries $1,560$
Max Depth $128$